the enthalpy of combustion of benzene is -3301Kj/mol a calorimeter has a heat capacity of 80j/c. how much benzene do you have to burn to heat the calorimeter up to 30c
Answer #1

Heat required by calorimeter,

Q = Ccal * delta T

= 80 J/oC * 30 oC

= 2400 J

= 2.400 KJ

mol of benzene required = -Q / heat of combustion

= - 2.400 KJ / -3301 KJ/mol

= 7.27*10^-4 mol

Molar mass of C6H6,

MM = 6*MM(C) + 6*MM(H)

= 6*12.01 + 6*1.008

= 78.108 g/mol

use:

mass of C6H6,

m = number of mol * molar mass

= 7.27*10^-4 mol * 78.108 g/mol

= 5.678*10^-2 g

Answer: 5.68*10^-2 g